A silicon monoxide thin film (n = 1.45) of thickness 83.2 nm is applied to a camera lens made of glass (n = 1.55). This will result in destructive interference for reflected light of what wavelength?

483 nm
386 nm
350 nm
603 nm
643 nm
Answer #1

Given

Refractive index of silicon monoxide μs = 1.45

Refractive index of glass μg = 1.55

Thickness of the film t = 83.2 nm

Solution

Here one light wave is reflected at the surface of the silicon dioxide top

One will go through this and then get get reflected by the glass and go through the refraction once again.

To give us the destructive interference the path difference must be λ/2

st = λ/2

2 x 1.45 x 83.2 x 10-9 = λ/2

λ = 482.56 x 10-9 m

λ = 483 nm